OpenStudy (anonymous):

You have a 3-card deck containing a king, a queen, and a jack. You draw a random card, then put it back and draw a second random card. Calculate the probability that you draw exactly 1 king.

OpenStudy (lgbasallote):

probability you draw a king at first is 1/3 then probability you DONT draw a king is 2/3 so simply multiply 1/3 x 2/3 then you have your answer

OpenStudy (zarkon):

KK KQ KJ QK QQ QJ JK JQ JJ 4 out of 9 have one king

OpenStudy (lgbasallote):

wait...so im wrong?

OpenStudy (zarkon):

P(one king) =P(1st king 2nd not king)+P(1st not king and 2nd is king) \[=\frac{1}{3}\times\frac{2}{3}+\frac{2}{3}\times\frac{1}{3}=\frac{4}{9}\]

OpenStudy (lgbasallote):

hmm i see